OpenStudy (anonymous):

2\sin t \cos t - \cos t + 2 \sin t -1 =0

OpenStudy (callisto):

Factorize the left side: \[2\sin t \cos t - \cos t + 2 \sin t -1 =0 \]\[(2\sin t -1)(1+\cos t) = 0\]Then put each factor = 0